(A) Given: C = 100 + 0.75Yd (where Yd = Y-T)

 I = 120-600i

 G = 200

 T = 20 + 0.2Y

 Ms/P = 300

Md/P = 50+0.5Y-600i

Where: C = Consumption

Y = Income

 I = Investment

G = Government spending

T = Taxes

i = interest rate

Ms/P = RealMoney Supply

Md/P = Real Demand for Money

(a) Derive the IS and LM curves

 

(b) Obtain the equilibrium:

i. Interest rate

ii. Income and consumption


1
Expert's answer
2021-05-17T10:38:11-0400

(a)

"Y = C+I+G"

"Y = 100 + 0.75(Y-20-0.2Y) + 120-600i + 200"

"Y = 420+0.6Y-15-600i"

"0.4Y = 405-600i"

"IS\space curve : Y = 1012.5 - 1500i"

 

Money Demand = Money Supply,

"50+0.5Y-600i = 300"

"LM\space curve : i =\frac{(0.5Y-250)}{600}"

 

(b)

(i)


"i =\frac{ (0.5Y-250)}{600}"

"i = 0.189"

Equilibrium interest rate : "i = 18.9\%"

(ii)

Substituting value of i in IS curve equation, we get,

Substituting value of i in IS curve equation, we get,

"Y = 1012.5 - 1500 \times \frac{ (0.5Y-250)}{600}"

"Y = 1012.5 - 1.25Y + 625"

"2.25Y = 1637.5"

Equilibrium Income : "Y = 727.7"

 

"C = 100+0.75(Y-20-0.2Y)"

"C = 100+0.6Y-15"

Equilibrium Consumption : "C = 521.6"